A small circular hole 5.00mm in diameter is cut in the side of a large water tank, 11.0 m below the water level in the tank. The water tank is sealed and contains air above the water at a gauge pressure of 2.00 atm.

a) What is the speed of efflux of the water coming out of the hole?

b) what is the volume discharged per second?


1
Expert's answer
2021-03-09T15:30:01-0500

a)


"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(11)}=14.7\frac{m}{s}"

b)


"C=14.7(0.25\pi(0.005))^2=2.27\cdot10^{-4}\frac{m^3}{s}"